x^2+y^3-x^2y^2=64 find x and y intercepts
Answer by lwsshak3(11628)

x^2+y^3-x^2y^2=64 find x and y intercepts
**
x-intercept:
set y=0
x^2=64
x=√64
x=±8
..
y=intercept:
set x=0
y^3=64
y=cube root of 64
y=4